9) Four times the difference of a number and one is equal to six times the sum of the number and three. Find the number.
Answer by edjones(7311)

x= the number
4(x-1)=6(x+3)
4x-4=6x+18
-22=2x
x=-11
Check:
4(-11-1)=6(-11+3)...-48=-48
Ed